OpenStudy (anonymous):

Julie throws a ball to her friend Sarah. The ball leaves Julie's hand a distance 1.5 meters above the ground with an initial speed of 16 m/s at an angle 47 degrees; with respect to the horizontal. Sarah catches the ball 1.5 meters above the ground. What is the maximum height the ball goes above the ground? I understand that the velocity should = 0, and I think I found the time to be 2.39 seconds, but when I plug everything in it always turns out wrong =( I use the formula d=vt-1/2at^2 since v should = 0 I do d=-1/2at^2 but when I plug the values in the answer is wrong. Please help!

OpenStudy (anonymous):

Consider horizontal component: Using \[v^2=u^2+2as\]\[0=(16\sin47^{0})^2+2(-9.81)s\]\[s=6.98m\]\[=7.0m(2s.f.)\] Hence, maximum height is 8.5m